OpenStudy (mylahb):

write in standard form: (-6-7i)+(-3-8i) @Aleah.wood0613

OpenStudy (michele_laino):

please, you have to eliminate theparentheses, so we ahve: \[-6-7i-3-8i=(-6-3)+i(-7-8)=...\] please complete the computation

OpenStudy (michele_laino):

what is -6-3=...?

OpenStudy (anonymous):

-9 - 15 i